OpenStudy (anonymous):

Calculating the Future Value of Property. Ben Collins plans to buy a house for $ 220,000. If that real estate is expected to increase in value by 3 percent each year, what will its approximate value be seven years from now?

OpenStudy (anonymous):

This is a question where you use an exponential equation. Exponential equations are always in the form y=a(b)^x a = initial value (when x = 0) b = growth rate So for this question your equation should look like this y=220000(1.03)^x Plug in 7 for x.

OpenStudy (anonymous):

Thank you